OpenStudy (anonymous):

Solve for x: 5 over quantity x squared minus 1 minus 2 over x equals 2 over quantity x plus 1

OpenStudy (campbell_st):

is this your equation \[\frac{5}{x^2 -1} - \frac{2}{x} = \frac{2}{x+1}\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok...get rid of the denominators by multiplying every term by x(x^2 -1) \[\frac{ 5\times x(x^2 -1)}{x^2 -1} - \frac{2\times x(x^2 -1)}{x} = \frac{2\times x(x^2-1)}{x+1}\] this is a little long.. sorry... but the right hand side can be simplified because x^2 -1 = (x+1)(x-1) so you end up with \[5x -2(x^2 -1) = 2x(x-1)\] simplify and collect like terms \[4x^2 -7x - 2 = 0\]

OpenStudy (campbell_st):

this factorises to (x - 2)(4x + 1) = 0 then x = 2 or - 1/4